# New Method for Robust Analysis of Critical Behavior of Magnetic Systems

### v1

@@ Abstract
Figures 2(a, b, c) show the analysis of the behavior of $\delta$, $\beta$ and $\gamma$ in different ranges of $T$ and $B$ as determined using our proposed method as described above for $T\leq T_{\rm{C}}$. In Figs. 2(b-i), seven different $T$ ranges of $10.7$% to $1.5\%$ of $T_{\rm{C}}$ are considered. We find that the CEs thus obtained are robust across different $T$ (for $T\leq T_{\rm{C}}$) and $B$ ranges clearly suggesting that the critical behavior associated with local electron moments (here, the modeled Ni system) is independent of $B$. The CE values thus determined correspond to the standard UC of 3D XY. We next proceed to determine the appropriate range of $T$ for $T>T_{\rm{C}}$ as follows. Figures 2 (d-f) show the representative plots of NSs, calculated from MAPs using the robust values of CEs obtained using the proposed method for $T\leq T_{\rm{C}}$. For brevity, three different $B$ ranges, viz., 5-15 T, 20-30 T and 30-40 T, respectively, are shown. Clearly, for $T\leq T_{\rm{C}}$, the $NS\sim 1$ for the entire $T$ range. However, for $T>T_{\rm{C}}$, we observe that the values deviate from $NS\sim 1$ beyond $T=T_{0}$ [Fig. 2(d)] which is very close to $T_{\rm{C}}$. Interestingly, the $T$-range, i.e., $T_{0}-T_{\rm{C}}$ for $T>T_{\rm{C}}$ showing $NS\sim 1$ grows with increasing $B$ [Figs. 2(e and f)]. We, therefore, propose a process to determine the appropriate ranges of $T$ and $B$ for critical analysis. Selecting the $M-H$ isotherms at relatively high-$B$ and low-$T$ ranges, $NSs$ are to be determined using the $\beta$ and corresponding $\gamma$ as we proposed above. The maximum $T$-range for the chosen $B$ range showing $NS\sim 1$ will be the appropriate range to be considered for the critical analysis. It is to be noted that the field-dependent critical analysis is required if $\delta$ varies with $B$. Furthermore, as we verify the behavior of linearity of $NSs$ with $T$ [Figs.2(g-i)], we find that the range of linearity of the intercepts also increases with increasing $B$-range confirming the identical critical behavior on either side of $T_{\rm{C}}$. We thus show that the appropriate $T$-range can also be determined from the plot of intercepts vs. $T$ showing the linear behavior.
@@
Figure 3: (a) The $M-T$ recorded for Gd in field-cooled (FC) and zero-field-cooled (ZFC) warming modes by applying $10$ mT of magnetic field along [0001]. Inset is the first-derivative of $M-T$ which yields $T_{\rm{C}}=292.4$ K. (b) The field-dependent variation of $\delta_{MH}$, estimated for different applied magnetic field ($B$) ranges as mentioned. (c) Variation of $T_{1}$ and $T_{2}$ which constitute the crossover temperature $T$ for the intercepts which are obtained from the linear fit to MAPs created using the $\beta$ and $\gamma$ of short-range 3D standard UCs and MF UC, respectively. (d) Variation of $NS_{\rm{avg}}$ vs. $\beta$ to estimate appropriate $\beta$ and $\gamma$ associated with Gd in the $B$-range from 4 to 5 T. (e) Variation of $\beta$ and $\gamma$, determined using our proposal, for different $B$ ranges. (f) $NS$ vs. $T$ plots, constructed using the CEs obtained using our proposal, for different $B$ ranges. SR: short-range, LR: long-range.
@@
Using our proposed methods for critical analysis, we next carry out to investigate the role of itinerant electron magnetism on the critical behavior emerging in Gd. Although the magnetic properties of this rare earth material is interesting for applications, however, the exact nature of the interactions leading to the observed magnetic behavior of Gd remains unclear [10, 39, 40, 39, 41, 42]. The complex competing interactions between local $4f$ as well as itinerant electrons makes Gd an interesting “nontrivial” system [40, 39, 41, 42]. The previous studies for critical behavior of Gd report conflicting conclusions [21, 22, 20, 23]. From $B$-dependent studies, it was suggested that the 3D Heisenberg UC in low-$B$ changes to 3D Ising in the high-$B$ [23]. However, the values of $\gamma$ and $\delta$ reported for low-$B$ range do not fully justify the claim. Intriguingly, although the role of long-range RKKY interaction involving itinerant electrons was suggested for Gd [10, 8], so far the analysis of critical behavior using existing methods indicated the presence of short-range interactions only [21, 22, 20, 23].
@@
We use high-quality single crystalline Gd in hexagonal close packed form, grown by Czochralski method [43], for the critical analysis [31]. The magnetization ($M-T$ and $M-H$) measurements were carried out in a SQUID setup (make: Quantum Design) by applying $B$ along [0001] [31]. The $T_{\rm{C}}$ for the crystal is 292.4 K [Fig. 3(a)], and the magnetic moment per Gd atom is found to be $\approx 7.63\,\mu_{B}$ which is consistent with the values reported in literature [44]. The unusual feature observed near $T_{C}$ in $M-T$ plot [Fig. 3(a)] is ascribed to the competing energies in the system [40, 39, 41, 42, 10]. Due to this, determination of $\beta$ using SP [Figs. 1(a) and 1(b)] is not straight forward. Therefore, we use our proposed methods for critical analysis as follows. Unlike for Ni [Fig. 2(a)], our results for Gd clearly show a variation of $\delta$ with $B$, see Fig. 3(b). The range of $\delta$ observed in this case covers nearly the values for long-range MF to that for short-range standard UCs. So, we proceed for $B$-dependent critical analysis by considering 8% of $T_{\rm{C}}$ as the appropriate $T$ range. The intercepts of the linear fit to MAPs, constructed using the CEs for short-range standard UCs as well as long-range MF UC, when plotted against the corresponding temperatures exhibit a crossover at $T_{1}$ and $T_{2}$, respectively [Fig. 3(c)] [31]. See Fig. 1(f) for corresponding plots for Ni. According to our proposal, the intercepts of MAPs obtained using CEs of either short-range UCs or long-range MF UC exhibiting a crossover at $T_{\rm{C}}$ would determine whether the system has a short-range or a long-range interaction, respectively. Thus, our analysis showing $T_{2}\sim T_{\rm{C}}$ at low-$B$ and deviating away from $T_{\rm{C}}$ at high-$B$ suggests that the long-range interactions emerging due to itinerant electron moments becomes weaker with $B$. On the other hand, $T_{1}$ starting at about $6$ K away from $T_{\rm{C}}$ [Fig. 3(c)] in the low-$B$ range remains almost constant in the entire measured range of $B$ suggesting that at high-$B$, the nature of interactions are predominantly of short-range. This highlights the $B$-dependence of magnetic interactions in Gd.
@@
Interestingly, independent analysis of CEs for Gd using our new approach shows results [Figs. 3(d) and 3(e)] which are consistent with that obtained from the intercept analyses. Whereas in low-$B$ range $\beta$ and $\delta$ suggest dominance of MF type long-range interaction, for high-$B$ the values suggest the interaction predominantly of 3D-Ising type short-range nature, see Figs. 3(b) and  3(e). Moreover, $\gamma$ is found to remain $\sim$ 1 in the entire measured range of $B$ [Fig. 3(e)] indicating the role of a complex competing interaction in Gd. Hence, our new approach based on the analysis of intercepts as well as CEs clearly identifies the competing role of short- and long-range interactions in different $B$-ranges in Gd. We argue that the local Gd moments interact via 3D Ising type short-range interaction while the presence of itinerant electrons results in MF type long-range RKKY interaction. We note that the CEs for typical magnetic systems, exhibiting competing interactions resulting in the unusual magnetic behavior, do not fall in a single UC [12, 6, 7, 13]. These observations eventually support our new findings. We again check the robustness and reliability of the obtained CE values (for $T\leq T_{\rm{C}}$) for Gd from the $NS$ plots which show $NS\sim 1$ in the entire $T\leq T_{\rm{C}}$ and in a small range of $T$ above $T_{\rm{C}}$, see Fig. 3(f). Similar to the case of Ni, the range of $T$ showing $NS\sim 1$ beyond $T_{\rm{C}}$ grows with $B$ indicating further that the critical behavior remains same on either side of $T_{\rm{C}}$.
@@
In conclusion, we propose novel approaches to determine two important parameters for the analysis of critical behavior of magnetic systems, viz., the range of interactions and the values of CEs. We demonstrate the successful application of the methods for two magnetic systems, viz. Ni and Gd. By analyzing the intercepts of MAPs, we appropriately identify the range of exchange interaction in a system. From computational and experimental investigations of the critical behavior of Ni and Gd, we claim that $NSs\sim 1$ for all $M-H$ isotherms for $T\leq T_{\rm{C}}$ is the necessary and sufficient condition to test the reliability and robustness of the values of the CEs as determined by our proposed method. We argue that the nontrivial magnetic systems with competing interactions may simultaneously exhibit two types of UCs resulting in effective values of CEs. Our investigations show that the strength of short-range interaction among local moments remains same while the RKKY type long-range interaction emerging in Gd, due to the involvement of itinerant electron moments, weakens with increasing $B$. These investigations thus solve the long-standing problem of the values of CEs not belonging to any specific UC observed for nontrivial systems. Our proposed methods may, thus, enable detailed understanding of critical behavior of magnetic systems.
@@ 1. Materials and Methods
Face centered cubic (FCC) Ni system was taken for the temperature- and field-dependent magnetization calculation. All the calculations were performed with the vampire software [45, 30] using the atomistic model simulations of magnetic Ni. The atomistic Landau–Lifshitz–Gilbert (LLG) equation was solved using the Heun integration scheme to calculate the magnetization curves [45]. The details about the LLG-Heun method can be found in [30, 45]. For the calculations for FCC Ni system, we take the following parameters: cluster size of 20$\times$20$\times$20 $nm$, unit cell size of 3.524 Å, atomistic spin moment $\mu_{S}=0.606$$\mu_{B}$, exchange energy $J_{ij}=2.757\times 10^{-26}$ J/link and we used magnetocrystalline anisotropy energy $k_{u}=0.0$ J/atom. For the LLG simulation, we use 50000 equilibrium steps and 500000 averaging steps. For all the calculations, we use the critical value of the Gilbert damping parameter equal to $1.0$ and the computational time step equal to $0.1$ fs [30, 45, 46]. We calculated temperature-dependent magnetization (M-T) with and without applying magnetic field. Finally, we converted the M-T isofields into field-dependent magnetization (M-H) isotherms and, thus, performed the theoretical investigation of the critical behavior of Ni system using the methods described below in sections 3 to 5.
@@
A single crystal (SC) of gadolinium (Gd) was grown using the Czochralski method [43] in a tetra-arc furnace (Technosearch Corporation, Tokyo, Japan). Approximately 7 grams of 99.9% pure Gd metal was placed on a water-cooled copper hearth inside the furnace. After evacuating the chamber to a vacuum of 10-6 mBar, it was filled with high-purity argon gas. The Gd metal was melted using the conventional arc melting process. After the initial melt, the Gd was flipped and remelted under the same argon atmosphere. Gadolinium melts at around 1300 °C. A previously prepared polycrystalline Gd seed crystal was carefully introduced into the molten Gd, causing local solidification around the seed. The seed crystal was then rapidly pulled at a speed of approximately 50 mm/h. The melt temperature was gradually increased to initiate the necking process, after which it was gradually decreased to widen the diameter of the growing crystal. Once steady-state conditions were achieved, the crystal was pulled at a steady speed of 10 mm/h for around 7 hours. The pulled ingot was subsequently aligned along its principal crystallographic directions using Laue diffraction and cut, using a wire EDM, along the basal plane and $c$-axis for anisotropic studies. Since the crystal was prone to oxidation in ambient conditions, precautions were taken to minimize its exposure to the atmosphere.
Temperature- and field-dependent magnetic measurements of SC Gd were carried out using Quantum Design magnetic properties measurement system (MPMS). The magnetic field was applied along [0001] direction.

### v2

@@ Abstract
Figures 2(a, b, c) show the analysis of the behavior of $\delta$, $\beta$ and $\gamma$ in different ranges of $T$ and $B$ as determined using our proposed method, as described above, for $T\leq T_{\rm{C}}$. In Figs. 2(b-i), seven different $T$ ranges, viz., $10.7$% of $T_{\rm{C}}$ to $1.5\%$ of $T_{\rm{C}}$ and three different ranges of $B$ are considered for all the relevant parameters discussed above. From Figs. 2(a-c) we find that the CEs thus obtained are robust across different $T$ (for $T\leq T_{\rm{C}}$) and $B$ ranges clearly suggesting that the critical behavior associated with local electron moments is independent of $B$. Figures 2 (d-f) show the representative plots of $NS$s, calculated from MAPs using these robust values of CEs obtained using the proposed method for $T\leq T_{\rm{C}}$. For brevity, three different $B$ ranges, viz., 5-15 T, 20-30 T and 30-40 T, respectively, are shown. Clearly, the $NS\sim 1$ for the entire $T$ range for $T\leq T_{\rm{C}}$. However, for $T>T_{\rm{C}}$, we observe that the values deviate from $NS\sim 1$ beyond $T=T_{0}$ [Fig. 2(d)] which is very close to $T_{\rm{C}}$. Interestingly for $T>T_{\rm{C}}$, the $T$-range, $T_{0}-T_{\rm{C}}$ showing $NS\sim 1$ grows with increasing $B$ [Figs. 2(e and f)]. We thus confirm that the CE values will remain the same on either side of $T_{\rm{C}}$ if one select the temperature range up to $T_{0}$ for $T>T_{\rm{C}}$. We, therefore, propose a process to determine the appropriate ranges of $T$ and $B$ for critical analysis. Selecting the $M-H$ isotherms at relatively high-$B$ and low-$T$ ranges, $NS$s are to be determined using the $\beta$ and corresponding $\gamma$ as we proposed above. The maximum $T$-range for the chosen $B$ range showing $NS\sim 1$ will be the appropriate range to be considered for the critical analysis. It is to be noted that the $B$-dependent critical analysis is required if $\delta$ varies with $B$. Furthermore, as we verify the behavior of linearity of $NS$ with $T$ [Figs.2(g-i)], we find that the range of linearity of the $I$s also increases with increasing $B$-range. We thus show that the appropriate $T$-range can also be determined from the plot of $I$s vs. $T$ exhibiting the linear behavior.
@@
Figure 3: (a) The $M-T$ recorded for Gd in field-cooled (FC) and zero-field-cooled (ZFC) warming modes by applying $B=10$ mT along [0001]. Inset is the first-derivative of $M-T$ which yields $T_{\rm{C}}=292.4$ K. (b) The $B$-dependent variation of $\delta_{MH}$, estimated for different $B$ ranges as mentioned. (c) Variation of the crossover temperatures $T_{1}$ and $T_{2}$ obtained from the $I$s of the linear fit to MAPs constructed using the CE values of the short-range standard UCs and MF UC, respectively. (d) Plots of $NS_{\rm{avg}}$ vs. $\beta$ to estimate appropriate $\beta$ and $\gamma$ associated with Gd in the $B$-range from 4 to 5 T. (e) Variation of $\beta$ and $\gamma$, determined using our proposal, for different $B$ ranges. (f) $NS$ vs. $T$ plots, constructed using the CEs obtained using our proposal, for different $B$ ranges. SR: short-range, LR: long-range.
@@
We next apply our proposed methods to investigate the role of itinerant electron magnetism on the critical behavior of Gd. Although the magnetic properties of this rare earth material is interesting for applications, however, the exact nature of the interactions leading to the observed magnetic behavior of Gd is complex and still remains unclear [10, 36, 37, 36, 38, 39]. The complex competing interactions between local $4f$ as well as itinerant electrons makes Gd an interesting “nontrivial” magnetic system [37, 36, 38, 39]. The previous studies for critical behavior of Gd report conflicting conclusions as regards to the nature of the magnetic interaction [22, 23, 21, 24]. Although the role of long-range Ruderman-Kittel-Kasuya-Yusida (RKKY) interaction involving itinerant electrons was suggested for Gd [10, 8], so far the analysis of critical behavior using existing methods indicated the presence of short-range interactions only [22, 23, 21, 24]. From $B$-dependent studies, it was also suggested that the 3D Heisenberg UC in low-$B$ changes to 3D Ising in the high-$B$ [24]. However, the values of $\gamma$ and $\delta$ reported for low-$B$ range do not fully justify the claim. To test our proposed method of critical analysis for this nontrivial magnetic system, we use high-quality single crystalline Gd in hexagonal close packed form, grown by Czochralski method [40, 30]. The magnetization ($M-T$ and $M-H$) measurements were carried out in a SQUID setup (make: Quantum Design) by applying $B$ along [0001] [30]. The $T_{\rm{C}}$ for the crystal is 292.4 K [Fig. 3(a)], and the magnetic moment per Gd atom is found to be $\approx 7.63\,\mu_{B}$ which is consistent with the values reported in literature [41]. The unusual characteristic features observed near $T_{C}$ in $M-T$ plot [see Fig. 3(a)] is ascribed to the competing energies in the system [37, 36, 38, 39, 10]. Due to this, determination of $\beta$ using SP [Figs. 1(a) and 1(b)] is not straight forward. This serves as a prerequisite for applying our proposed methods to analyze the critical behavior of Gd.
@@
Unlike for Ni [Fig. 2(a)], our results for Gd clearly show a variation of $\delta$ with $B$, see Fig. 3(b). The range of $\delta$ observed in this case covers nearly the values for long-range MF to that for short-range standard UCs. So, we proceed for $B$-dependent critical analysis by considering 8% of $T_{\rm{C}}$ as the appropriate $T$-range for $T\leq T_{\rm{C}}$. The $I$s of the linear fit to MAPs, constructed using the CEs for short-range standard UCs as well as long-range MF UC, when plotted against the corresponding temperatures exhibit a crossover at $T_{1}$ and $T_{2}$, respectively [30]. For reference, see Fig. 1(f) for the corresponding plots for Ni. According to our proposal discussed above, the $I$s of MAPs obtained using CEs of either short-range UCs or long-range MF UC exhibiting a crossover at $T_{\rm{C}}$ would determine whether the system has a short-range or a long-range interaction, respectively. Our analysis, which shows that $T_{2}\sim T_{\rm{C}}$ at low-$B$ but deviates from $T_{\rm{C}}$ at high-$B$ [Fig. 3(c)], suggests that the long-range interaction mediated by itinerant electron moments weakens with increasing $B$. The Lorentz force acts to localize itinerant electrons, thereby suppressing their long-range interaction characteristics [42, 43]. On the other hand, $T_{1}$ starting at about $6$ K away from $T_{\rm{C}}$ [Fig. 3(c)] in the low-$B$ range remains almost constant in the entire measured range of $B$ suggesting no effect of $B$ on the range of exchange interaction associated with local electron moments. This competing effect indicates that at high $B$, the interactions are predominantly short-range in nature due to the weakening of the RKKY interaction.
@@
Next, we determine CE values associated with Gd using our new approach, see Fig. 3(d) [30]. Here, $B$-dependent critical analysis is prerequisite as value of $\delta$ changes with $B$. The CE values determined for Gd show results [Fig. 3(e)] which are consistent with that obtained from the $I$s [Fig. 3(c)] analyses. The $B$-dependent effective values of $\beta$ and $\delta$ indicate the dominance of MF type long-range interaction in low-$B$ and predominantly 3D Ising type short-range interaction at high $B$, see Figs. 3(b) and  3(e). Moreover, $\gamma$ increases monotonically and found to remain close to 1 in the entire measured range of $B$ [Fig. 3(e)] indicating the role of a complex competing interactions in Gd. One should note here that the critical analysis yields the effective values of CEs. Interestingly, this has gone largely unnoticed. It remains unclear why the range and type (Heisenberg, Ising, XY, etc.) of exchange interactions would vary with $B$ in a simple ferromagnet. While conventional analyses [25] indicate predominantly long-range interactions, supported by $\sigma\approx 1.5$ [30] across the entire field range, our intercept analysis reveals a clear competing effect of short- and long-range interactions with increasing $B$ in Gd. Based on our investigations using the proposed new approaches, we argue that the local Gd moments interact via 3D Ising-type short-range interaction, while the presence of itinerant electrons gives rise to MF-type long-range RKKY interaction. The competition gives rise to effective critical exponent (CE) values that vary with $B$, reflecting the coexistence of multiple UCs arising from distinct exchange interactions. With increasing $B$, the itinerant electrons begin to localize [42], resulting in the suppression of long-range interactions, while the local electron moments remain unaffected, leading to a reduction in the $\beta$ value. Our investigations in this work thus clarifies the reasons behind the CE values, for nontrivial systems, not belonging to any single UC [11, 6, 7, 12]. Thus, it can be argued that the critical behavior of a magnetic system heavily depends on the local environment around the magnetic ion(s) [18, 20, 6, 12]. If the local environment of the magnetic ions changes with the field, the CE values will also vary with the field, and vice versa. We again check the robustness and reliability of the obtained CE values (for $T\leq T_{\rm{C}}$) for Gd from the $NS$ plots which show $NS\sim 1$ in the entire $T\leq T_{\rm{C}}$ and in a small range of $T$ above $T_{\rm{C}}$, see Fig. 3(f). Similar to the case of Ni, the range of $T$, showing $NS\sim 1$ beyond $T_{\rm{C}}$, grows with $B$ indicating further that the critical behavior remains same on either side of $T_{\rm{C}}$.
@@
In conclusion, we propose new methods to determine two important parameters for the analysis of critical behavior of magnetic systems, viz., the range of magnetic interactions and the values of CEs. We demonstrate the successful application of the methods for two magnetic systems, viz. Ni and Gd. By analyzing the $I$s of MAPs, we appropriately identify the range of exchange interaction in a system. From computational and experimental investigations of the critical behavior of Ni and Gd, we claim that $NSs\sim 1$ for all $M-H$ isotherms for $T\leq T_{\rm{C}}$ is the necessary and sufficient condition to test the reliability and robustness of the CE values as determined using our proposed method. We argue that the magnetic systems with competing interactions may simultaneously exhibit two types of UCs resulting in effective values of CEs. Our investigations show that the strength of short-range interaction among local moments remains same while the RKKY type long-range interaction emerging in Gd, due to the involvement of itinerant electron moments, weakens with increasing $B$. These investigations thus solve the long-standing problem of the values of CEs not belonging to any specific UC. The methods and approaches discussed in this article is equally applicable for all types of magnetic systems holding the necessary conditions required for critical analysis. Our proposed methods may thus enable a detailed understanding of the critical behavior of magnetic systems.
@@ 1. Materials and Methods
Face centered cubic (FCC) Ni system was taken for the temperature- and field-dependent magnetization calculation. All the calculations were performed with the vampire software [31, 29] using the atomistic model simulations of magnetic Ni. The atomistic Landau–Lifshitz–Gilbert (LLG) equation was solved using the Heun integration scheme to calculate the magnetization curves [31]. The details about the LLG-Heun method can be found in [29, 31]. For the calculations for FCC Ni system, we take the following parameters: cluster size of 20$\times$20$\times$20 $nm$, unit cell size of 3.524 Å, atomistic spin moment $\mu_{S}=0.606$$\mu_{B}$, exchange energy $J_{ij}=2.757\times 10^{-26}$ J/link and we used magnetocrystalline anisotropy energy $k_{u}=0.0$ J/atom. For the LLG simulation, we use 50000 equilibrium steps and 500000 averaging steps. For all the calculations, we use the critical value of the Gilbert damping parameter equal to $1.0$ and the computational time step equal to $0.1$ fs [29, 31, 32]. We calculated temperature-dependent magnetization (M-T) with and without applying magnetic field. Finally, we converted the M-T isofields into field-dependent magnetization (M-H) isotherms and, thus, performed the theoretical investigation of the critical behavior of Ni system using the methods described below in sections $\mathrm{III}$ to $\mathrm{V}$.
@@
A single crystal (SC) of gadolinium (Gd) was grown using the Czochralski method [40] in a tetra-arc furnace (Technosearch Corporation, Tokyo, Japan). Approximately 7 grams of 99.9% pure Gd metal was placed on a water-cooled copper hearth inside the furnace. After evacuating the chamber to a vacuum of 10-6 mBar, it was filled with high-purity argon gas. The Gd metal was melted using the conventional arc melting process. After the initial melt, the Gd was flipped and remelted under the same argon atmosphere. Gd melts at around 1300 °C. A previously prepared polycrystalline Gd seed crystal was carefully introduced into the molten Gd, causing local solidification around the seed. The seed crystal was then rapidly pulled at a speed of approximately 50 mm/h. The melt temperature was gradually increased to initiate the necking process, after which it was gradually decreased to widen the diameter of the growing crystal. Once steady-state conditions were achieved, the crystal was pulled at a steady speed of 10 mm/h for around 7 hours. The pulled ingot was subsequently aligned along its principal crystallographic directions using Laue diffraction and cut, using a wire EDM, along the basal plane and $c$-axis for anisotropic studies. Since the crystal was prone to oxidation in ambient conditions, precautions were taken to minimize its exposure to the atmosphere.
Temperature- and field-dependent magnetic (M-T and M-H) measurements of SC Gd were carried out using Quantum Design magnetic properties measurement system (MPMS). The magnetic field was applied along [0001] direction.